OpenStudy (anonymous):

Identify Chemical Reaction Type 2 Ag + S = Ag2S Ca(OH)2+2HCl=CaCl2+2H 2 C4H10 + 13 O2 = 8 CO2 + 10 2 KClO3 = 2 KCl + 3 O2 2 AgNO3 + Cu = Cu(NO3)2 + 2

OpenStudy (chmvijay):

2 Ag + S = Ag2S redox reaction 2 C4H10 + 13 O2 = 8 CO2 + 10 combustion 2 KClO3 = 2 KCl + 3 O2 decomposition 2 AgNO3 + Cu = Cu(NO3)2 + 2Ag displacement reaction

OpenStudy (chmvijay):

Ca(OH)2+2HCl=CaCl2+2H is neutralization reaction or acid base reaction

OpenStudy (anonymous):

Thank you!